OpenStudy (anonymous):

^3sqrt(2x-1) differentiate

OpenStudy (anonymous):

differentiate y = log base e 3x+8x^3

sam (.sam.):

\[\Large \frac{d}{dx}(\log_e(3x+8x^3))\]

sam (.sam.):

Is it?

sam (.sam.):

\[y = \log_e (3x+8x^3)\] Using \(\LARGE\log_ab=c \Leftarrow\Rightarrow b=a^c\), gives \[\Large 3x+8x^2=e^y\] Can you differentiate it now? it's implicit differentiation.

OpenStudy (raden):



OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, the derivative of cube root of f(x) is f ' (x)/ 3 cube root of (f(x))^2 so,